In a market for dry cleaning, the inverse market demand function is given by P = 250 - Q and the (private) marginal cost of production for the aggregation of all dry-cleaning firms is given by MC = 50 + Q. Finally, the pollution generated by the dry-cleaning process creates external damages given by the marginal external cost curve MEC = Q + 5.

1) What will be the equilibrium output produced in a perfectly competitive industry?

a. 150

b. 130

c. 120

d. 100

2) What will be the market price in a perfectly competitive industry?

a. $180

b. $160

c. $150

d. $120

3)What is the efficient level of output that should be produced in this market?

a. 100

b. 85

c. 65

d. 50

4) What should be the efficient market price that will take into account the external cost of pollution to society?

a. $125

b. $145

c. $165

d. $185
Answer #1

1)

P=250-Q

MC=50+Q

In case of perfectly competitive market, Set P=MC for profit maximization

250-Q=50+Q

200=2Q

Q=100 (equilibrium output)

Correct option is

d) 100

2)

Given P=250-Q

Put Q=100

P=250-100=$150 (equilibrium price)

Correct option is

c) $150

3)

Marginal Social Cost=MSC=MC+MEC=50+Q+Q+5=55+2Q

Set MSC=P for socially optimal output

55+2Q=250-Q

3Q=195

Q=65 (socially optimal output)

Correct option is

c) 65

4)

Given P=250-Q

Put Q=65

P=250-65=$185 (socially optimal price)

Correct option is

d) $185